What is the probability of at least one head?
If you throw a single fair coin multiple times, the probability of getting NO head is:For 1 throw: 1/2
For 2 throws: 1/2 squared = 1/4
For 3 throws: 1/2 cubed = 1/8
etc.
The probability of getting AT LEAST ONE head is the complement; for example, for 3 throws, it would be 1 minus 1/8.
What is the probability that the spinner and number cube both land on 2?
That would depend on how many numbers are on the spinner and the cube. The more numbers there are, the less likely it is that they would both land an any given number.
What does probability measures?
From a scale of 1 to 0 probability measure an event will happen, maybe it will happen or it will not happen.
In the Empirical Rule of data will fall in with two standard deviation.?
Approx 95% of the observations.
If you flip a coin what is the chance of getting tails?
A coin has two (major) sides and neither side has any more advantage than the other.
100 (percent) divided by 2 (sides) = 50
There is a 50% chance (or 1/2) of the coin landing on tails.
What is a probability that when a couple has three children at least one of them is a boy?
There are two problems in answering this question. The gender of children are not independent events and also, the probability of a boy is not half as is often assumed but nearer to 52% and it varies over time and between countries.
However, if you assume that the genders are independent and that the probability of either gender is 0.5 then the answer is 0.875
4/13
As the two events are independent multiply the respective probabilities together:
pr(5) = 1/6
pr(tails) = 1/2*
→ pr(5 and tails) = 1/6 × 1/2 = 1/12
*Actually it is slightly less than 1/2 as there is a very small probability that the coin will end up on its edge; for a small coin it is so small that it can be ignored, but with a scaled up coin it becomes a real possibility as (accidentally) demonstrated in the 1997 Royal Institution Christmas Lecture "The Magical Maze"
one in three
1 2 = under 3
3 4 5 6 = at least 3
so 4/6 probability
coin is 1 in 2
4/6 x 1/2 = 4/12 = 1/3
Pr(Both picked 3) = 1/16Pr(Both picked the same) = 1/4.
What is The set of all possible outcomes in a probability experiment is the?
It is the outcome space.
How many '6 of hearts' are in a standard 52 card deck?
One. There is one card for each combination of number and suite.
Think, if there are A, E, I, O, U, and couting Y, there are 6 vowels. there are 26 letters. So the answer is 6 over 26 and then you simplify it to 3 over 13.
If you flipped a coin 60 times how many times would you expect the coin to land on heads?
30 maybe but i say 35 or 31
about 23.9
The theoretical probability of randomly picking each color marble is the number of color marbles you have for each color, divided by the total number of marbles. For example, the probability of selecting a red marble is 3/20.
Provided the random variable can take only finitely many, equally spaced values, you can write one in terms of the other. For details see the paper by Ales Cerny entitled "Introduction to Fast Fourier Transform in finance".
With replacement, the odds of the second extraction do not change after the first, because of independence of the two extractions. Hence, the probability of extracting two reds is (3/5)*(3/5), the probability of two greens is (2/5)*(2/5) and that of one red and one green is 2*(3/5)*(2/5).
What is the probability that the sequence will be Heads-Tails-Heads if you toss a coin 3 times?
Since there are 2 outcomes for a coin toss, and you will toss the coin 3 times the number of outcomes are 23 or 8. Since H-T-H can occur only 1 way, the probability of the H-T-H sequence is 1/8.
Why is strata called homogenous in stratified sampling?
Homogeneous refers to groups composed of parts or elements that are all of the same kind or nature. In stratified sampling, a population which is composed of diverse groupings is subdivided into two or more groups so that the diversity is decreased in the subgroups. For example, if the total population is composed of males and females, then stratification into subgroups of male and female will result in strata that are of the same kind with respect to the classification variable gender: i.e, the strata are homogeneous. Other classification variables or combinations of classification variables may be used to improve homogeneity.
Why is the mean the standard partner of the standard deviation?
The mean and standard deviation often go together because they both describe different but complementary things about a distribution of data. The mean can tell you where the center of the distribution is and the standard deviation can tell you how much the data is spread around the mean.
IF YOU HAVE ONE GAME, YOU HAVE A POSIBILITY OF 3 OUTCOMES, IF YOU HAVE 2 GAMES IT GO'S UP TO 9 POSIBILE OUTCOMES.
FIRST ANSWER 1X3=3
SECOND IS 3X3=9
FOR 10 GAMES IT IS 3X3 TEN TIMES
3X3X3X3X3X3X3X3X3X3=59,049
POSSIBLE COMBINATIONS
(I THINK)HA HA
For each question the student has a 50% chance of answering right.
We can express this as a fraction of 1 (1 being 100%): 1/2, i.e. 0.5
There are 5 questions, so the answer will be:
0.5*0.5*0.5*0.5*0.5=0.5^5=0.03125
or 3.125%